6. UK’s real GDP was 1 360 trillion pounds [£] in 2009 and 1 434 trillion pounds in 2010. UK’s population was 191.5 million people in 2009 and 193.3 million in 2010.
Calculate.
a. the economic growth rate.[3]
b. The growth rate of real GDP per person [5]
c. The approximate number of years it takes for real GDP per person in the UK to double if the 2010 economic growth rate and population growth rate are maintained.[4]
1
Expert's answer
2019-08-02T10:32:29-0400

a. "\frac{1434-1360}{1360}" =0.05441


b.The growth rate of real GDP per person= "\frac{7.419-7.102}{7.102}" = 0.0446


c. The GDP per person = 4.46% it will double= using the rule of 72 = "\frac{72}{4.46}" = 16 year